Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn, Page 2 of 4  Not logged in ELOG logo
ID Date Author Group Subject Text Attachments
  49   12 Apr 2025, 23:15 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 500 uJ, 246 pulses, stable
delivery
  
  48   11 Apr 2025, 23:14 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 246 pulses,
stable delivery
  
  47   10 Apr 2025, 23:17 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 128 pulses, stable
delivery 
  
  46   09 Apr 2025, 22:46 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV, ~ 400 uJ, 128 pulses, stable
delivery 
  
  45   09 Apr 2025, 10:22 Jan GruenertPRCSA3 statusSASE3 Apple-X undulator cell 26
issue - SOLVED

During the night, an issue appeared
  
  44   09 Apr 2025, 09:41 Jan GruenertPRCIssueSASE1 XTD2 attenuator arm issue

On monday 7.4.2025, one of the
SASE1 attenuator arms (SA1_XTD2_ATT/PNACT/A9)
  
  43   07 Apr 2025, 15:35 Jan GruenertPRCStability IssueSASE1 pulse energy drops slowly
three times

SASE1
XGM detects three times during the morning
 JG_2025-04-07_um_15.34.41.png 
  42   07 Apr 2025, 07:19 Harald SinnPRCSA1 status7:15 SPB reports large background from
non-lasing SASE3 bunches. BKR is informed.  

7:50 BKR analysed situation. Everything
  
  41   06 Apr 2025, 06:33 Harald SinnPRCSA2 status5:49 HED called and they have problems
with the Laser Sync signal. A change of laser
delay is not possibe. We supsect that it
 LBsync1.jpgLBsync2.jpgLBsync3.jpgLBsync4.jpgLBsync5.jpg 
  40   03 Apr 2025, 07:34 Harald SinnPRCSA1 status7:20 FXE reports a problem with laser synchronisation
(yellow warning light and could not change
delay). BKR informs the DESY laser sync group. 
  
  39   27 Mar 2025, 22:57 Tokushi SatoSPBShift end**X-ray delivery:**
* 16 keV, ~upto 1 mJ, 160 pulses
**Optical Laser:**
  
  38   17 Mar 2025, 06:51 Romain LetrunSPBIssueThe status of the FXE shutter in the Public
status display DOOCS panel is incorrect.
Here appears as opened, when it is actually
 Screenshot_from_2025-03-17_06-45-52.png 
  37   15 Mar 2025, 19:49 Harald SinnPRCStatusMID called and reported issue with Dachs
terminals at the user labs (probably no power).
TS was informed. Because the lab are
  
  36   12 Mar 2025, 23:14 Raphael de WijnSPBShift summary
SPB/SFX
shift summary user experiment 6933 (setup
day)
  
  35   04 Mar 2025, 07:05 Tokushi SatoSPBShift summaryX-ray delivery

    up to 188/352
pulses per train (~1 mJ pulse energy)@13.4
  
  34   03 Mar 2025, 07:05 Tokushi SatoSPBShift summaryX-ray delivery

    up to 188 pulses
per train (~1 mJ pulse energy)@13.4 keV,
  
  33   02 Mar 2025, 06:56 Tokushi SatoSPBShift summaryX-ray delivery

    up to 175 pulses
per train (~1 mJ pulse energy)@13.4 keV,
  
  32   01 Mar 2025, 11:53 Martin DommachPRCIssueDue to issues
SCS has with pulses from SASE1 lasing in
  
  31   01 Mar 2025, 06:56 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202/352
pulses per train (~0.8 mJ pulse energy)@13.4
  
  30   28 Feb 2025, 06:57 Tokushi SatoSPBShift summaryX-ray delivery

    up to 202/352
pulses per train (~0.8 mJ pulse energy)@13.4
  
ELOG V3.1.4-7c3fd00